Circular No. 7260

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET 1999 S2 R. H. McNaught, Australian National University, reports his discovery

of a comet on an R survey film taken with the U.K. Schmidt Telescope by F. G. Watson. The comet shows a very strong central condensation, a weak circular coma of diameter 20′′, and a diffuse tail 3′.5 long in p.a. 210^{o}. Confirming CCD images by McNaught with the 1.0-m f/8 reflector at Siding Spring on Sept. 21 yield m_{2} = 20.1–20.2.

1999 UT α_{2000} δ_{2000} m_{1} Observer
Sept.19.72016 3^{h}13^{m}21.^{s}68 −63^{o}48′28′′.8 17.5 Watson
19.76183 3 13 20.56 −63 48 33.5  _{”}
21.55226 3 12 26.29 −63 53 46.0  McNaught
21.55392 3 12 26.24 −63 53 46.2  _{”}
21.55595 3 12 26.16 −63 53 46.6  _{”}
21.55789 3 12 26.09 −63 53 46.9  _{”}

DO DRACONIS This variable is in outburst for the first time since mid-Oct. 1996 (cf.

IAUC 6493), when it reached m_{v} = 10.4, according to J. A. Mattei, AAVSO. The following visual magnitude estimates were forwarded by E. O. Waagen, AAVSO: Sept. 18.81 UT, [13.5 (P. Schmeer, Bischmisheim, Germany); 20.830, 11.2 (E. Muyllaert, Oostende, Belgium); 20.904, 11.0 (M. Westlund, Uppsala, Sweden); 21.142, 10.7 (Schmeer).

V4444 SAGITTARII D. K. Lynch, R. J. Rudy, S. Mazuk, The Aerospace Corporation; and

R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego, report 0.8–2.5-μm spectrophotometry of V4444 Sgr = N Sgr 1999 using the Shane 3-m telescope (+ Aerospace Near Infrared Imaging Spectrograph) at Lick Observatory on Aug. 31.19 UT (∼ 120 days after outburst). The spectrum shows many emission lines (FWHM ∼ 2000 km/s) with the strongest being He I 1.0830-μm, followed by the Paschen and Brackett lines. Also present were the Lyβ fluoresced O I lines at 0.8446 and 1.1287 μm and the [Si VI] coronal line at 1.9629 μm. The spectrum also showed a continuum rising smoothly to the red between 1.5 and 2.4 μm. This could indicate thermal emission from a dust shell or the effects of strong interstellar reddening. J, H, and K magnitudes were 12.4, 12.0, and 10.0, respectively.

1999 September 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green